[ The area of a right triangle of which the base is 3 times the height, is 216 cm squared. calculate the base of the triangle. ]
Answer by TimothyLamb(4379)

a = (1/2)bh
---
b = 3h
h = (1/3)b
---
a = (1/2)bh
216 = (1/2)b(1/3)b
216 = (1/2)(1/3)bb
216 = (1/6)bb
bb = 1296
---
Answer:
b = 36 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php